A number is multiplied by 4. Then 7 is added to the product. After subtractg 3, the result is 8. What is the number
Answer by ReadingBoosters(3246)

x = the number
(4x + 7) - 3 = 8
4x + 4 = 8
4x + 4 - 4 = 8 - 4
4x = 4
4x/4 = 4/4
x = 1
Proof
(4)(1) = 4 + 7 = 11 - 3 = 8